# AXP4, a Numerical Simulator of Pulse Profiles for Binary Accreting X-ray Pulsars – I: Beam Geometries

### v1

@@ 1 Introduction
Since the magnetic pressure is far stronger than the radiation and gas pressure, it is the dominant factor in governing X-ray propagation. In the presence of a strong magnetic field around the neutron star, the outgoing X-radiation gets collimated along a preferential direction of propagation, due to magneto-scattering with the accreted plasma matter. The result is an emission ‘beam’ emanating from the polar caps. As a pulsar rotates about an axis misaligned with the magnetic dipole axis, its emission beams sweep across the observer’s line-of-sight, producing pulsating measurements of X-ray flux for an observer on Earth, as though a galactic X-ray ‘light-house’. When the anisotropic beams sweep across the observer’s line-of-sight, the emission is detected as periodic, coherent ‘pulses’ from the target source in the X-ray band. The observed X-ray pulse profiles are investigated in the paper improving previous attempts. Several of these are simulated for various source configurations by the custom, generalized ‘Accreting X-ray Pulsar Pulse Profile Producer’ (AXP4) code developed in this work.
@@
Work in this paper is proposed to be devoted to a final model and detailed investigation of beamed emission in accreting binary X-ray pulsars and building a custom tool to predict these observables for a wide range of physical conditions at high resolution to obtain updated, refined estimates of the initial numerical models for high mass X-ray binaries. Observables for different geometries, defined by the relative angles between the spin axis, the magnetic axis and the view axis can be computed. The resulting simulated light curves can be directly compared with observations from the existing and upcoming missions (see accompanying paper).
@@ 1.1.3 Homogeneous atmosphere
In Section 5, the simulations are further improved by developing another new module, ‘Composite Slab Integrator’ which breaks Meszaros & Nagel’s assumption of point emission regions and accounts for a more realistic geometry of hotspots by integrating the emission over a circular slab at the polar caps. Such comprehensive composite gravitationally bent and hotspot-integrated pulses are also simulated at high resolution. For the sake of theoretical completeness, this study is further extended in the accompanying paper to examine the effect of an increasing slab area (to geometrically spherical caps on the neutron star) on the pulsed fraction, for the full span of surface coverage (0%–100%) at high resolution with a new, custom ‘Surface Coverage’ module (cylindrical surfaces can be constructed for accretion columns). Modified pulse profiles for circular slabs are thus obtained by integrating the emission over two antipodal circular slabs lying flush on the neutron star surface at the antipodal polar caps, with Meszaros & Nagel’s point emission functions used for modeling the emission from each infinitesimal surface area element. The two modules so developed can be run individually or in conjunction with each other. The resultant model composite surface-integrated and relativistically modified profile simulations are reported in Figs. 9. For comparison, we supplement these profiles generated using model emission beams with alternative pulse profiles generated using hypothetical, fiducial input beams from uniform, isotropic test emitters in Section 6.6.
@@ 4.1 Photon trajectories in curved space-time
This is a straight-forward but approximate equation used to relate the angles of initial launching and the final ray asymptote towards the observer through the strength of the gravitational field i.e. the curvature of traversed space-time. It is valid for trajectories beyond two Schwarzschild radii () from the center of the gravitational potential well. This holds well for binary accreting X-ray pulsars, where the emission occurs from the polar caps . If the error is defined as with , then Beloborodov’s approximation is accurate to an error of within which goes to up to for the maximum value of (Kumar, 2012).
@@ 8 Future Scope
Such a cornerstone code presented in this paper for circular slabs could be extended step-wise to accommodate other sub-geometries like multiple ovals (Miller et al., 2019; Riley et al., 2019) using suitable limits of integration for an elliptical grid and an additive operation for the net emission, crescents (Miller et al., 2019) using a subtractive operation on individual emission from two sets of antipodal hotspots with an inter-slab offset, rings (full, fractional and partial-filled) using suitable limits of integration for the polar grid on the slab, including multiple-temperature components with a suitable modification in the beam input and for non-antipodal (including both in the same rotational sphere) hotspots with unshared parameters (Miller et al., 2019) not excluding the contribution from the second hotspot. The exercise with gravitational light bending and red-shift can be extended to accretion columns as per the prescription by De Falco et al. (2016), accounting for the gravitational gradient along the height of the column. The computed results can be compared with observations obtained from various global X-ray satellites – both existing and upcoming. The accompanying paper compares these simulations with Cen X-3 observations analyzed in Shirke et al. (2021).
The work carried out in this thesis and the subsequent results presented are polarisation-summed. Given the onset of global X-ray polarimetric space missions opening a new window of high energy polarisation and ushering in an era of X-ray polarisation studies, these results can be resolved into the and polarisation modes maintaining a separate treatment for each mode. To this end, a description of how the prescription used by Meszaros et al. (1988) can be customized for the work is outlined in detail in (Shirke, 2024). Elements from polarised radiative transfer (Mościbrodzka & Gammie, 2018) and parallel transport and bending of polarised light (Bronzwaer et al., 2020) would have to be additionally incorporated. Since polarimetric studies aim to resolve the pencil and fan beam degeneracy, alternate theoretical approaches for addressing this issue through the compactness parameter, are being considered.
The AXP4 simulator is intended to be made publicly available (open access) after adding some user-friendly layers and along with the requisite documentation121212Subject to academic tenure.

### v2

@@ 1 Introduction
Since the magnetic pressure is far stronger than the radiation and gas pressure, it is the dominant factor in governing the direction of X-ray propagation. In the presence of a strong magnetic field around the neutron star, the outgoing X-radiation gets collimated along a preferential direction of propagation, due to magneto-scattering with the accreted plasma matter. The result is an emission ‘beam’ emanating from the polar caps. As a pulsar rotates about an axis misaligned with the magnetic dipole axis, its emission beams sweep across the observer’s line-of-sight, producing pulsating measurements of X-ray flux for an observer on Earth, as though a galactic X-ray ‘light-house’. When the anisotropic beams sweep across the observer’s line-of-sight, the emission is detected as periodic, coherent ‘pulses’ from the target source in the X-ray band. The observed X-ray pulse profiles are investigated in the paper improving previous attempts. Several of these are simulated for various source configurations by the custom, generalized ‘Accreting X-ray Pulsar Pulse Profile Producer’ (AXP4) code developed in this work.
@@
Work in this paper is proposed to be devoted to a final model and detailed investigation of beamed emission in accreting binary X-ray pulsars and building a custom tool to predict these observables for a wide range of physical conditions at high resolution to obtain updated, refined estimates of the initial numerical models for high mass X-ray binaries. Observables for different geometries, defined by the relative angles between the spin axis, the magnetic axis and the view axis can be computed. The resulting simulated light curves can be directly compared with observations from the existing and upcoming missions (see accompanying paper, Shirke & Dewangan (2024)).
@@ 1.1.3 Homogeneous atmosphere
In Section 5, the simulations are further improved by developing another new module, ‘Composite Slab Integrator’ which breaks Meszaros & Nagel’s assumption of point emission regions and accounts for a more realistic geometry of hotspots by integrating the emission over a circular slab at the polar caps. Such comprehensive composite gravitationally bent and hotspot-integrated pulses are also simulated at high resolution. For the sake of theoretical completeness, this study is further extended in the accompanying paper, Shirke & Dewangan (2024) to examine the effect of an increasing slab area (to geometrically spherical caps on the neutron star) on the pulsed fraction, for the full span of surface coverage (0%–100%) at high resolution with a new, custom ‘Surface Coverage’ module (cylindrical surfaces can be constructed for accretion columns). Modified pulse profiles for circular slabs are thus obtained by integrating the emission over two antipodal circular slabs lying flush on the neutron star surface at the antipodal polar caps, with Meszaros & Nagel’s point emission functions used for modeling the emission from each infinitesimal surface area element. The two modules so developed can be run individually or in conjunction with each other. The resultant model composite surface-integrated and relativistically modified profile simulations are reported in Figs. 9. For comparison, we supplement these profiles generated using model emission beams with alternative pulse profiles generated using hypothetical, fiducial input beams from uniform, isotropic test emitters in Section 6.6.
@@ 4.1 Photon trajectories in curved space-time
This is a straight-forward but approximate equation used to relate the angles of initial launching and the final ray asymptote towards the observer through the strength of the gravitational field i.e. the curvature of traversed space-time. It is valid for trajectories beyond two Schwarzschild radii () from the center of the gravitational potential well. This holds well for binary accreting X-ray pulsars, where the emission occurs from the polar caps (See Appendix B). If the error is defined as with , then Beloborodov’s approximation is accurate to an error of within which goes to up to for the maximum value of (Kumar, 2012).
@@ 8 Future Scope
Such a cornerstone code presented in this paper for circular slabs could be extended step-wise to accommodate other sub-geometries like multiple ovals (Miller et al., 2019; Riley et al., 2019) using suitable limits of integration for an elliptical grid and an additive operation for the net emission, crescents (Miller et al., 2019) using a subtractive operation on individual emission from two sets of antipodal hotspots with an inter-slab offset, rings (full, fractional and partial-filled) using suitable limits of integration for the polar grid on the slab, including multiple-temperature components with a suitable modification in the beam input and for non-antipodal (including both in the same rotational sphere) hotspots with unshared parameters (Miller et al., 2019) not excluding the contribution from the second hotspot. The exercise with gravitational light bending and red-shift can be extended to accretion columns as per the prescription by De Falco et al. (2016), accounting for the gravitational gradient along the height of the column. The computed results can be compared with observations obtained from various global X-ray satellites – both existing and upcoming. The accompanying paper, Shirke & Dewangan (2024) compares these simulations with Cen X-3 observations analyzed in Shirke et al. (2021).
The work carried out in this thesis and the subsequent results presented are polarisation-summed. Given the onset of global X-ray polarimetric space missions opening a new window of high energy polarisation and ushering in an era of X-ray polarisation studies, these results can be resolved into the and polarisation modes maintaining a separate treatment for each mode. To this end, a description of how the prescription used by Meszaros et al. (1988) can be customized for the work is outlined in detail in Shirke (2024). Elements from polarised radiative transfer (Mościbrodzka & Gammie, 2018) and parallel transport and bending of polarised light (Bronzwaer et al., 2020) would have to be additionally incorporated. Since polarimetric studies aim to resolve the pencil and fan beam degeneracy, alternate theoretical approaches for addressing this issue through the compactness parameter, are being considered.
The AXP4 simulator has been beta-tested and is intended to be made publicly available (open access) after adding some user-friendly layers and along with the requisite documentation121212Subject to academic tenure. As seen in Fig. 13, the radius of the neutron star has been further parametrized in the same.
@@ Appendix A High-resolution profiles
Appendix B Exact light bending
The exact formula for the gravitational bending of light has been provided by Misner et al. (1973) which relates and in a Schwarzschild geometry as,
where is the impact parameter,
Fig. 13a shows a comparison between the implementation of Beloborodov’s approximate gravitational light bending with the exact formula (See Eqs. (A1) and (A2) of Poutanen & Beloborodov (2006)) for an AXP4 simulated pulse profile. As expected from the justification provided in Sec. 4.1, the two results are coincident.
Appendix C Special relativistic correction
Fig. 13b displays the skewed Doppler boosted profiles for two antipodal isotropic slab emitters of a fast rotating pulsar. The bolometric flux in Eq. (A16) of Poutanen & Beloborodov (2006) is reduced to in the absence of gravitational bending and surface integration (a typical spectral power law index of is chosen for accreting millisecond pulsars). This figure can be compared with their Fig. 2 which provides results for only a single hotspot.